Assume an overall game standard deviation of 0.94 per unit and casino advantage of 1.15% (approximate value for a large series of bets equally divided between player and banker). Calculate the 95% confidence limits for the total casino win in a series of 1000 hands at $300 per hand. O (-$22381, +$34881) O (-$4676,+$6976) O (-$9352,+$13952) O (-$14028, +$20929) O None of the above.
Answer #1

here for total expected amount=1000*300*1.15%=3450

and std deviation =0.94*300*sqrt(1000)=8917.623

for 95% CI ; crtiical z =1.96

therefore 95% confidence interval for total casino win=3450-/+1.96*8917.623= -14028.0 ; 20929